OpenStudy (anonymous):

Using only 2's and any combination of mathematical signs,symbols,and functions, can you make 5 ?

OpenStudy (anonymous):

2*2=4*2=8+2=10/2=5

OpenStudy (anonymous):

That's one way to make 5